# Controllability score for linear time-invariant systems on an infinite time horizon

### v1

@@ Abstract
We introduce a scaled controllability Gramian that can be computed reliably even for unstable systems. Using this scaled Gramian, we reformulate the controllability scoring problems into equivalent but numerically stable optimization problems. Their optimal solutions define dynamics-aware network centrality measures, referred to as the volumetric controllability score (VCS) and the average energy controllability score (AECS). We then formulate controllability scoring problems on an infinite time horizon. Under suitable assumptions, we prove that the resulting VCS and AECS are unique and that the finite-horizon scores converge to them. We further show that VCS and AECS can differ markedly in this limit, because VCS enforces controllability of the full system, whereas AECS accounts only for the stable modes. Finally, using Laplacian dynamics as a representative example, we present numerical experiments that illustrate this convergence.
@@ I Introduction
The controllability scores are designed to address limitations of existing controllability-based frameworks. Prior approaches can be broadly categorized into structural controllability [15] and quantitative controllability measures [18]. Structural controllability depends only on topology and enables efficient graph-theoretic analysis without identifying edge weights [16, 17], but it does not preclude prohibitively large control-energy requirements [30]. Quantitative measures incorporate edge weights and time horizons to capture energy-related aspects of controllability; building on this viewpoint, controllability scores evaluate node importance while accounting for control effort. Existing Gramian-based centralities such as volumetric controllability centrality (VCE) and average control energy centrality (ACE) may still be misaligned with quantitative controllability because they rely on the rank or pseudoinverse of a single-input controllability Gramian [23]. By contrast, controllability scores are defined by introducing virtual actuation at all state nodes and optimizing the input allocation, rather than relying on a single-input controllability Gramian. Subsequent studies established uniqueness properties [23, 22], demonstrated applicability to human brain networks [22], and incorporated constraints on actuated nodes [24].
In Section II-A, we define two controllability scores (VCS and AECS) as optimal solutions to distinct controllability-Gramian–based optimization problems. This finite-horizon formulation requires specifying a terminal time $T>0$. However, the choice of $T$ has not been systematically addressed in the existing literature. It has been reported in [22] that when $T$ is small, the resulting scores are dominated by short-term dynamics and fail to capture meaningful system-wide properties, whereas for sufficiently large $T$, they reflect the network-structured dynamics of the system. These observations suggest taking $T$ large. A natural extreme is $T=\infty$, which eliminates dependence on the system time scale and on the researcher’s subjective choice, and is therefore desirable from the viewpoint of reproducibility. However, except for the case where $A$ is Hurwitz (i.e., all its eigenvalues have strictly negative real parts), the controllability Gramian diverges as $T\to\infty$, so the scores are no longer well defined. Moreover, even for large but finite $T$, the Gramian can become ill-conditioned, making objective-function evaluation numerically unstable and potentially preventing accurate computation. This paper provides a solution to this issue.
@@
    By taking the limit $T\to\infty$ in the reformulated problems, we obtain controllability scores on an infinite time horizon, and prove, under suitable assumptions, that the finite-horizon scores converge to them. This provides a principled choice of the terminal time, namely $T=\infty$, and also offers an explanation for the empirical observation reported in [22] that, although the controllability Gramian diverges for unstable systems, the controllability scores nevertheless converge to a certain value as $T\to\infty$. This infinite-horizon formulation also highlights an important advantage of defining centrality through optimization problems. While several centrality measures based on functions of the controllability Gramian have been proposed [25], including trace- or log-determinant–based ones, these quantities are generally not well-defined on an infinite time horizon unless the system matrix $A$ is Hurwitz. This is because the controllability Gramian diverges as $T\to\infty$, as discussed in Remark 1. In contrast, since the controllability scores are defined as the solutions to optimization problems over a bounded feasible set, the scores themselves can remain meaningful even when the controllability Gramian diverges.
@@
    Under suitable assumptions, we show that the scaled controllability Gramian can be computed in a numerically stable manner using a real Schur decomposition and the solution of a Sylvester equation. This enables the stable numerical evaluation of the infinite-horizon controllability scores.
@@
Throughout this paper, $n$ denotes the number of nodes in the network, and $A\in\mathbb{R}^{n\times n}$ denotes the system matrix in 1. The sets of all natural numbers, all real numbers, and all complex numbers are denoted by $\mathbb{N}$, $\mathbb{R}$, and $\mathbb{C}$, respectively. Let $e_{i}\coloneq(0,\ldots,0,1,0,\ldots,0)^{\top}\in\mathbb{R}^{n}$ be a standard vector whose $i$-th entry is $1$ and other entries are 0. Let $I_{k}$ and $N_{k}$ denote the identity matrix of size $k$ and the nilpotent Jordan block of size $k$ with ones on the superdiagonal, respectively. Let $O$ denote a zero matrix of a suitable size.
Let $\Delta$ and $\widetilde{\Delta}$ be the standard simplices in $\mathbb{R}^{n}$ and $\mathbb{R}^{n-1}$, respectively, i.e.,
$\Delta\coloneq\left\{(p_{i})\in\mathbb{R}^{n}\,\middle|\,\begin{gathered}p_{i}\geq 0\ (i=1,\ldots,n),\\ \textstyle\sum_{i=1}^{n}p_{i}=1\end{gathered}\right\},$
The mapping $\varphi\colon\widetilde{\Delta}\to\Delta$ defined by
is a homeomorphism; its inverse is given by
$\varphi^{-1}(p_{1},\ldots,p_{n})=(p_{1},\ldots,p_{n-1})^{\top}.$
For $\lambda\in\mathbb{C}$, $\operatorname{Re}(\lambda)$, $\operatorname{Im}(\lambda)$, and $\overline{\lambda}$ denote the real part, the imaginary part, and the complex conjugate of $\lambda$, respectively. For a real matrix $\in\mathbb{R}^{k\times\ell}, XX^{\top}$ denotes the transpose of $X$, and for a complex matrix $\in\mathbb{C}^{k\times\ell}, YY^{*}$ denotes the conjugate transpose of $Y$. For a real invertible matrix $X\in\mathbb{R}^{\ell\times\ell}$, we write $X^{-\top}\coloneq(X^{-1})^{\top}$, and for a complex invertible matrix $Y\in\mathbb{R}^{\ell\times\ell}$, we write $Y^{-*}\coloneq(Y^{-1})^{*}$. For a matrix $Z\in\mathbb{C}^{\ell\times\ell}$, $\mathrm{e}^{Z},\det Z$, and $\operatorname{tr}Z$ denote the exponential of $Z$, the determinant of $Z$, and the trace of $Z$. For $C=(c_{ij})\in\mathbb{R}^{\ell\times\ell}$ and $^{\prime}\in\mathbb{R}^{\ell^{\prime}\times\ell^{\prime}}, CC$$\otimes C^{\prime}\coloneq\begin{pmatrix}c_{11}C^{\prime}&\cdots&c_{1\ell}C^{\prime}\\ \vdots&\ddots&\vdots\\ c_{\ell 1}C^{\prime}&\cdots&c_{\ell\ell}C^{\prime}\end{pmatrix}\in\mathbb{R}^{\ell\ell^{\prime}\times\ell\ell^{\prime}}$ denotes their Kronecker product. For Hermitian matrices $X,Y\in\mathbb{C}^{\ell\times\ell}$, we write $X\succ Y$ if XYX-Y is positive definite. For $v_{1},\ldots,v_{\ell}\in\mathbb{C}$, let $\mathrm{diag}(v_{1},\ldots,v_{\ell})$ denote the diagonal matrix with the diagonal elements $v_{1},\ldots,v_{\ell}$. For a vector $v=(v_{1},\ldots,v_{\ell})^{\top}\in\mathbb{C}^{\ell}$, $\left\|{v}\right\|\coloneq\sqrt{\sum_{i=1}^{\ell}|v_{i}|^{2}}$ denotes the standard Euclidean norm. For a matrix $X\in\mathbb{C}^{k\times\ell}$, $\left\|{X}\right\|\coloneq\sqrt{\operatorname{tr}(X^{*}X)}$ denotes the Frobenius norm.
@@ II-A Controllability score
In this subsection, we summarize the concept of the controllability scores [23]. To define them, we introduce a virtual input allocation in which a fixed amount of control resources is distributed across network nodes. Let $p\in\Delta$ and consider the following system illustrated in Fig. 1:
@@
where $u(t)\coloneq(u_{1}(t),\ldots,u_{n}(t))^{\top}\in\mathbb{R}^{n}$ is a virtual control signal, and each node $x_{i}$ is actuated by $\sqrt{p_{i}}u_{i}(t)$. The value $p_{i}$ represents the amount of control resources allocated to node $x_{i}$; under a constraint on the input energy, a larger value of $p_{i}$ allows a stronger control effect on node $x_{i}$, indicating that node $x_{i}$ is a central node in the system. The controllability score associates each coefficient $p_{i}$ with the importance of node $x_{i}$.
@@
Figure 1: Virtual system for defining controllability scores. The node $x_{i}$ denotes a state variable, and the node $u_{i}$ denotes a virtual control signal. The coefficients $p_{i}$ specify how a fixed amount of control resources is allocated to the nodes, and each input $u_{i}$ corresponds one-to-one with node $x_{i}$ through the weight $\sqrt{p_{i}}$.
Let us maximize the controllability of system 2 with respect to $p$ for a certain measure, and let $p^{*}$ denote the optimal solution. The optimal solution $p^{*}$ represents the allocation of control authority that best enhances system controllability. The entries $p_{i}^{*}$ of this allocation, therefore, quantify how important each node $x_{i}$ is for controlling the system dynamics. Accordingly, $p^{*}$ can be regarded as a centrality measure, and we refer to it as a controllability score.
The controllability of a system can be measured by the controllability Gramian. For $B\in\mathbb{C}^{n\times n}$, let $\widetilde{W}(B;T)\in\mathbb{C}^{n\times n}$ be defined as follows:
@@
The controllability Gramian for 2 with terminal time $T>0$ is given by $\widetilde{W}(\operatorname{diag}(p);T)$, which we abbreviate as $\widetilde{W}(p;T)$. The quantities $\log\det\widetilde{W}(p;T)$ and $\operatorname{tr}\left(\widetilde{W}(p;T)^{-1}\right)$ can serve as scalar controllability measures. A smaller value of $\log\det\widetilde{W}(p;T)$ corresponds to a larger volume of the reachable set $\mathcal{E}(p;T)$, defined as
@@ II-B Remaining challenges and our approach
Note that centrality measures based on functions of the controllability Gramian are not necessarily well defined in the limit $T\to\infty$. For example, average controllability centrality (AC), VCE, and ACE proposed in [25] were originally defined for stable systems, but can be naively extended to unstable systems, in which case they are defined for node $i$ by
@@
where $k_{i}\coloneq\operatorname{rank}\left(\widetilde{W}(e_{i}e_{i}^{\top};T)\right)$ and $\lambda_{j}\left(\widetilde{W}(e_{i}e_{i}^{\top};T)\right)\ (j=1,\ldots,k_{i})$ denote the positive eigenvalues. Since eigenvalues of $\widetilde{W}(e_{i}e_{i}^{\top};T)$ that are close to zero reduce the values of $\mathrm{VCE}_{i,T}$ and $\mathrm{ACE}_{i,T}$, whereas zero eigenvalues do not, it has been pointed out in [23] that VCE and ACE do not necessarily provide appropriate centrality measures from the perspective of controllability.
Furthermore, as discussed in [22], capturing the influence of network structure requires choosing the terminal time $T$ to be sufficiently large, which naturally leads one to consider the limit $T\to\infty$. Since $\left\|{\widetilde{W}(e_{i}e_{i}^{\top};T)}\right\|\to\infty$ as $T\to\infty$ for at least one $i$, the maximum eigenvalue of $\widetilde{W}(e_{i}e_{i}^{\top};T)$ also diverges, implying $\mathrm{AC}_{i,T}\to\infty$ and $\mathrm{VCE}_{i,T}\to\infty$.
@@ III Scaling method of the controllability Gramian and reformulation of Problems 1 and 2
In this section, we reformulate Problem 1 as
@@ III-A Problem settings
In this section, the matrix $A\in\mathbb{R}^{n\times n}$ is arbitrary, and we assume that it admits a Jordan decomposition
@@
We further introduce additional notation for later use. Let $n_{-}\coloneq\sum_{i=1}^{m_{-}}n_{i},\ n_{0}\coloneq\sum_{i=m_{-}+1}^{m_{-}+m_{0}}n_{i}$, and $n_{+}\coloneq\sum_{i=m_{-}+m_{0}+1}^{m}n_{i}$ denote the numbers of eigenvalues with negative, zero, and positive real parts, respectively, counted with algebraic multiplicity. Let
$J_{-}\coloneq\begin{pmatrix}J_{1}\\ &\ddots\\ &&J_{m_{-}}\end{pmatrix}\in\mathbb{C}^{n_{-}\times n_{-}}$
be the Jordan matrix corresponding to the eigenvalues with negative real parts. Similarly, we define $J_{0}\in\mathbb{C}^{n_{0}\times n_{0}}$ and $J_{+}\in\mathbb{C}^{n_{+}\times n_{+}}$.

### v2

# Infinite-horizon controllability scores for linear time-invariant systems

@@ Abstract
We introduce a numerically stable reformulation of controllability scoring based on a scaled controllability Gramian, which remains reliably computable even for unstable systems. The resulting optimization problems define dynamics-aware network centrality measures, referred to as the volumetric controllability score (VCS) and the average energy controllability score (AECS). Building on this stable reformulation, we derive the corresponding infinite-horizon problems, develop an algorithm to solve them, and highlight computational advantages over their finite-horizon counterparts. Under suitable assumptions, we prove that the infinite-horizon VCS and AECS are unique and that the finite-horizon scores converge to them. We further show that VCS and AECS can differ markedly in this limit, because VCS enforces controllability of the full system, whereas AECS accounts only for the stable modes. Finally, numerical experiments on Laplacian dynamics illustrate this convergence.
@@ I Introduction
The controllability scores address limitations of existing controllability-based frameworks, which can be broadly categorized into structural controllability [16] and quantitative controllability measures [19]. Structural controllability depends only on network topology and enables graph-theoretic analysis without identifying edge weights [17, 18], but it does not preclude prohibitively large control-energy requirements [31]. Quantitative measures incorporate edge weights and time horizons to capture energy-related aspects of controllability; in this spirit, controllability scores quantify node importance while accounting for control effort. Existing Gramian-based centralities such as VCE and ACE introduced in [26] may still be misaligned with quantitative controllability because they rely on the rank or pseudoinverse of a single-input controllability Gramian [24]. By contrast, controllability scores are defined by introducing virtual actuation at all state nodes and optimizing the input allocation. Subsequent studies established uniqueness properties [24, 23], demonstrated applicability to human brain networks [23], and incorporated constraints on actuated nodes [25].
In Section II-A, we define two controllability scores, VCS and AECS, as optimal solutions to distinct controllability-Gramian–based optimization problems on a finite horizon. This formulation requires choosing a terminal time $T>0$. While existing work has not systematically addressed this choice, [23] reports that small $T$ yields scores dominated by short-term dynamics, whereas sufficiently large $T$ reveals network-structured, system-wide behavior, suggesting that $T$ should be large. A natural extreme is $T=\infty$, which removes the need to tune $T$ and improves reproducibility. However, unless $A$ is Hurwitz (i.e., all its eigenvalues have strictly negative real parts), the controllability Gramian diverges as $T\to\infty$, making the scores ill-defined. Moreover, even for large but finite $T$, the Gramian can be ill-conditioned, leading to numerically unstable evaluation of the objective and inaccurate computation. This paper addresses these issues.
@@
    By taking the limit $T\to\infty$ in the reformulated problems, we obtain controllability scores on an infinite time horizon, and prove, under suitable assumptions, that these scores are unique and that the finite-horizon scores converge to them. This provides a principled choice of the terminal time, namely $T=\infty$, and also offers an explanation for the empirical observation reported in [23] that, although the controllability Gramian diverges for unstable systems, the controllability scores nevertheless converge to a certain value as $T\to\infty$. This infinite-horizon formulation also highlights an important advantage of defining centrality through optimization problems. While several centrality measures based on functions of the controllability Gramian have been proposed [26], including trace- or log-determinant–based ones, these quantities are generally not well-defined on an infinite time horizon unless the system matrix $A$ is Hurwitz. This is because the controllability Gramian diverges as $T\to\infty$, as discussed in Remark 1. In contrast, since the controllability scores are defined as the solutions to optimization problems over a bounded feasible set, the scores themselves can remain meaningful even when the controllability Gramian diverges.
@@
    Under suitable assumptions, the infinite-horizon scaled controllability Gramian—and hence the infinite-horizon controllability scores—admits a numerically stable computation using the real Schur decomposition, followed by solving Sylvester and Lyapunov equations. Unlike the finite-horizon case, which typically requires numerical integration with a computational cost growing with $T$, the infinite-horizon formulation relies on solving Lyapunov equations with a cost independent of $T$. Moreover, this computation can be further accelerated using efficient approximation algorithms such as CF-ADI [15], making the infinite-horizon approach computationally advantageous.
@@
Throughout this paper, $n$ denotes the number of nodes in the network, and $A\in\mathbb{R}^{n\times n}$ denotes the system matrix in 1. The sets of all natural numbers, all real numbers, and all complex numbers are denoted by $\mathbb{N}$, $\mathbb{R}$, and $\mathbb{C}$, respectively. The symbol $\sqrt{-1}$ denotes the imaginary unit. Let $e_{i}\coloneq(0,\ldots,0,1,0,\ldots,0)^{\top}\in\mathbb{R}^{n}$ be a standard vector whose $i$-th entry is $1$ and other entries are 0. Let $I_{k}$ and $N_{k}$ denote the identity matrix of size $k$ and the nilpotent Jordan block of size $k$ with ones on the superdiagonal, respectively. Let $O$ denote a zero matrix of a suitable size. Let $\Delta$ be the standard simplex in $\mathbb{R}^{n}$, i.e.,
$\Delta\coloneq\left\{(p_{i})\in\mathbb{R}^{n}\,\middle|\,\begin{gathered}p_{i}\geq 0\ (i=1,\ldots,n),\\ \textstyle\sum_{i=1}^{n}p_{i}=1\end{gathered}\right\}.$
For $\lambda\in\mathbb{C}$, $\operatorname{Re}(\lambda)$, $\operatorname{Im}(\lambda)$, and $\overline{\lambda}$ denote the real part, the imaginary part, and the complex conjugate of $\lambda$, respectively. For a real matrix $\in\mathbb{R}^{k\times\ell}, XX^{\top}$ denotes the transpose of $X$, and for a complex matrix $\in\mathbb{C}^{k\times\ell}, YY^{*}$ denotes the conjugate transpose of $Y$. For a real invertible matrix $X\in\mathbb{R}^{\ell\times\ell}$, we write $X^{-\top}\coloneq(X^{-1})^{\top}$, and for a complex invertible matrix $Y\in\mathbb{R}^{\ell\times\ell}$, we write $Y^{-*}\coloneq(Y^{-1})^{*}$. For $Z\in\mathbb{C}^{\ell\times\ell}$, $\mathrm{e}^{Z},\det Z$, and $\operatorname{tr}Z$ denote the exponential, the determinant, and the trace of $Z$, respectively. For $C=(c_{ij})\in\mathbb{C}^{k\times\ell}$ and $^{\prime}\in\mathbb{C}^{k^{\prime}\times\ell^{\prime}}, CC$$\otimes C^{\prime}\coloneq\begin{pmatrix}c_{11}C^{\prime}&\cdots&c_{1\ell}C^{\prime}\\ \vdots&\ddots&\vdots\\ c_{k1}C^{\prime}&\cdots&c_{k\ell}C^{\prime}\end{pmatrix}\in\mathbb{C}^{kk^{\prime}\times\ell\ell^{\prime}}$ denotes their Kronecker product. For Hermitian matrices $X,Y\in\mathbb{C}^{\ell\times\ell}$, we write $X\succ Y$ if XYX-Y is positive definite. For $v_{1},\ldots,v_{\ell}\in\mathbb{C}$, let $\mathrm{diag}(v_{1},\ldots,v_{\ell})$ denote the diagonal matrix with the diagonal elements $v_{1},\ldots,v_{\ell}$. For $v=(v_{1},\ldots,v_{\ell})^{\top}\in\mathbb{C}^{\ell}$, $\left\|{v}\right\|\coloneq\sqrt{\sum_{i=1}^{\ell}|v_{i}|^{2}}$ denotes the standard Euclidean norm, and for $X\in\mathbb{C}^{k\times\ell}$, $\left\|{X}\right\|\coloneq\sqrt{\operatorname{tr}(X^{*}X)}$ denotes the Frobenius norm.
@@ II-A Controllability score
We begin with an intuitive description of controllability scores [24] to motivate the formal definitions introduced later in this subsection. To this end, we consider the autonomous system (1) and introduce a virtual actuation mechanism (see Fig. 1). Specifically, each state node $x_{i}$ is paired one-to-one with a virtual input $u_{i}$, and the input strength is determined by an allocation vector $p\in\Delta$:
@@
where $u(t)\coloneq(u_{1}(t),\ldots,u_{n}(t))^{\top}\in\mathbb{R}^{n}$. The value $p_{i}$ represents the amount of control resources allocated to node $x_{i}$; under a constraint on the input energy, a larger value of $p_{i}$ allows a stronger control effect on node $x_{i}$, indicating that node $x_{i}$ is a central node in the system. Given a controllability measure for system 2, we optimize it over $p$ and denote an optimal allocation by $p^{*}$. This optimal allocation $p^{*}$ maximizes controllability, and its entries $p_{i}^{*}$ quantify node importance. We therefore interpret $p^{*}$ as a dynamics-aware centrality measure, called a controllability score.
@@
Figure 1: Virtual system for defining controllability scores. Each node $x_{i}$ is paired one-to-one with a virtual input node $u_{i}$; the allocation $p$ distributes a fixed control budget, and $u_{i}$ acts on $x_{i}$ with gain $\sqrt{p_{i}}$.
To turn this intuition into a formal definition, we introduce controllability-Gramian–based scalar measures of controllability as functions of $p$. We first define, for $B\in\mathbb{C}^{n\times n}$ and $T>0$,
@@
The controllability Gramian for system 2 with terminal time $T>0$ is given by $\widetilde{W}(\operatorname{diag}(p);T)$, which we abbreviate as $\widetilde{W}(p;T)$. The quantities $\log\det\widetilde{W}(p;T)$ and $\operatorname{tr}\left(\widetilde{W}(p;T)^{-1}\right)$ can serve as scalar controllability measures. A smaller value of $\log\det\widetilde{W}(p;T)$ corresponds to a larger volume of the reachable set $\mathcal{E}(p;T)$, defined as
@@ II-B Remaining challenges and our approach
Controllability-Gramian–based centralities need not have a meaningful limit as $T\to\infty$. For example, average controllability centrality (AC), VCE, and ACE proposed in [26] (originally for stable systems) can be formally extended to unstable systems by defining, for node $i$,
@@
where $k_{i}\coloneq\operatorname{rank}\left(\widetilde{W}(e_{i}e_{i}^{\top};T)\right)$ and $\lambda_{j}\left(\widetilde{W}(e_{i}e_{i}^{\top};T)\right)\ (j=1,\ldots,k_{i})$ denote the positive eigenvalues. To capture the influence of network structure, one typically needs to take $T$ sufficiently large [23], which naturally motivates considering the limit $T\to\infty$. For unstable systems, however, one has $\left\|{\widetilde{W}(e_{i}e_{i}^{\top};T)}\right\|\to\infty$ as $T\to\infty$ for at least one $i$; hence the largest eigenvalue of $\widetilde{W}(e_{i}e_{i}^{\top};T)$ diverges, implying $\mathrm{AC}_{i,T}\to\infty$ and $\mathrm{VCE}_{i,T}\to\infty$. Even for finite $T$, VCE and ACE can be misleading as controllability-based centralities. In fact, as pointed out in [24], directions with very small (but positive) Gramian eigenvalues are heavily penalized, whereas directions with zero eigenvalues are omitted from the definitions and hence not penalized.
@@ III Scaling method of the controllability Gramian and reformulation of Problems 1 and 2
To define the infinite-horizon controllability scores in Section IV, we first introduce a scaling of the controllability Gramian and derive numerically stable reformulations of the finite-horizon problems Problems 1 and 2. Specifically, in this section we reformulate Problem 1 as
@@ III-A Problem settings
Throughout this section, the matrix $A\in\mathbb{R}^{n\times n}$ is arbitrary, and we assume that it admits a Jordan decomposition
@@
We further introduce additional notation for later use. Let $n_{-}\coloneq\sum_{i=1}^{m_{-}}n_{i},\ n_{0}\coloneq\sum_{i=m_{-}+1}^{m_{-}+m_{0}}n_{i}$, and $n_{+}\coloneq\sum_{i=m_{-}+m_{0}+1}^{m}n_{i}$ denote the numbers of eigenvalues with negative, zero, and positive real parts, respectively, counted with algebraic multiplicity. Let $J_{-}\coloneq\begin{pmatrix}J_{1}\\ &\ddots\\ &&J_{m_{-}}\end{pmatrix}\in\mathbb{C}^{n_{-}\times n_{-}}$ be the Jordan matrix corresponding to the eigenvalues with negative real parts. Similarly, we define $J_{0}\in\mathbb{C}^{n_{0}\times n_{0}}$ and $J_{+}\in\mathbb{C}^{n_{+}\times n_{+}}$.